OpenStudy (anonymous):

find all real solutions of the equation: (2x^4)+(4x^2)+1=0

OpenStudy (lgbasallote):

let a = x^2 so a^2 = x^4 so your equation will become something like 2a^2 + 4a + 1 = 0 now you can use quadratic formula to solve for a (because this is not factorable)

OpenStudy (lgbasallote):

then whatever value you get for a, take the square root to get the value of the solutions

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

Yes, that helps. Thank You.

OpenStudy (lgbasallote):

welcome